If the volume of a sphere is 12pie, what is the radius?
Answer by jim_thompson5910(35256)

V = (4/3)*pi*r^3

12pi = (4/3)*pi*r^3

12pi/pi = ((4/3)*pi*r^3)/pi

12 = (4/3)r^3

12*3 = 4r^3

36 = 4r^3

36/4 = r^3

9 = r^3

r^3 = 9

r = cube root of 9

r = 2.0800838 (this is approximate)